Every few years, winds in Boulder, Colorado, attain sustained speeds of 45.0 m/s (about 100 miles per hour) when the jet stream descends during early spring. Approximately what is the force due to the Bernoulli effect on a roof having an area of 245 m2? Typical air density in Boulder is 1.14 kg/m3, and the corresponding atmospheric pressure is 8.89 ✕ 104 N/m2. (Bernoulli's principle as stated in the text assumes laminar flow. Using the principle here produces only an approximate result, because there is significant turbulence.)
Answer #1

Here,

p, density of air = 1.14 Kg/m^3

for the force due to Bernoulli's effect

for the pressure difference , delta P = 0.50 * p * v1^2 - 0.50 * p * 0^2

delta P = 0.50 * 1.14 * 45^2

delta P = 1154.2 Pa

net force on the roof = delta P * area

net force on the roof = 1154.2 * 245

net force on the roof = 282791 N

the net force on the roof is 282791 N